Electrical engineering car battery. We measure the voltage across its terminals with a voltmeter and it shows an (open circuit) voltage of 12 volts. We place a 5 ohm resistor across the battery terminals, and we measure the current through with ammeter, and it shows 3 Amps. Show the Norton and Thevenin equivalent circuits including parameters for this battery.
Answer #1

place a votimeter 12V - 12V Here (v = voe=toy Here after connecting 52 Resistor 350 12V I- 3 A 12 and I= R +5 3= 12. R15 R +5be negativen Rth=8 The venius In 2 Nortons ar = 12A